The length of a retangle is twice the width. The area is 392 yards^2. Find the length and width.
I tried to solve this problem, but I am not sure if I did it correctly.
A=196(95)-(95)
=18620 - 9025
=9595
w=95
l=95

Answer by checkley77(12844)

DON'T KNOW WHERE YOU GET 30? BUT IT ISN'T IN THE PROBLEM NOR IN THE ANSWER.
THE ANSWER IS THE LENGTH IS 28 YDS & THE WIDTH 14 YDS. THIS GIVES AN AREA=392.
LW=AREA
L=2W
2W*W=392
2W^2-392=0
2(W^2-196)=0
2(W+14)(W-14)=0
W-14=0
W=14 ANSWER FOR THE WIDTH.
L=2*14=28 FOR THE LENGTH.
PROOF:
28*14=392
392=392